OpenStudy (anonymous):

A regular hexagon has sides of 5 feet. What is the area of the hexagon . if I did it right I got 12.5 root 3

OpenStudy (johnweldon1993):

Well let see \[\large A = \frac{3\sqrt{3}}{2}s^2\] s = 5 so we have \[\large \frac{3\sqrt{3}}{2}25\] \[\large \frac{75\sqrt{3}}{2}\] \[\large A = 37.5\sqrt{3}\]